How do I find, three consecutive odd numbers such that the sum of the first and second numbers is 23 less than three times the third?
Answer by CubeyThePenguin(3113)

consecutive integers: (x-2), x, (x+2)

(x-2) + x = 3(x+2) - 23
2x - 2 = 3x - 17
-x = -15
x = 15

The integers are 13, 15, and 17.